# Generalization Bounds via Meta-Learned Model Representations: PAC-Bayes and Sample Compression Hypernetworks

### v2

@@ 3.1 PAC-Bayes Encoder-Decoder Hypernetworks
We depart from previous works on PAC-Bayesian meta-learning and from classical PAC-Bayes, which consider a hierarchy of (meta-)prior and (meta-)posterior distributions (e.g., Pentina & Lampert, 2014). Instead, we consider distributions over a latent representation learned by the hypernetwork $\operatorname{\mathscr{H}}_{\theta}$. To encourage this representation to focus on relevant information, we adopt an encoder-decoder architecture $\operatorname{\mathscr{H}}_{\theta}(\cdot)=\operatorname{\mathscr{D}}_{\psi}(\operatorname{\mathscr{E}}_{\phi}(\cdot))$, with $\theta=\{\phi,\psi\}$. The encoder $\operatorname{\mathscr{E}}_{\phi}$ compresses the relevant dataset information into a vector ${\boldsymbol{\mu}}\in{\mathbb{R}}^{|{\boldsymbol{\mu}}|}$ (typically, $|{\boldsymbol{\mu}}|\ll|\gamma|$). The vector ${\boldsymbol{\mu}}$ is then treated as the mean of the Gaussian posterior distribution $Q_{\boldsymbol{\mu}}={\mathcal{N}}({\boldsymbol{\mu}},{\mathbf{I}})$ over the latent representation space, such that the decoder $\operatorname{\mathscr{D}}_{\psi}$ maps the realizations drawn from $Q_{\boldsymbol{\mu}}$ to the downstream parameters $\gamma$.
Training objective. Based on the above formulation, the training objective of the proposed encoder-decoder hypernetwork is
$\min_{\psi,\phi}\left\{\frac{1}{n}\sum_{i=1}^{n}\operatorname*{\mathbb{E}}\widehat{\mathcal{L}}_{\hat{T}_{i}}(h_{\gamma_{i}})\,\Bigg{|}\,\gamma_{i}{=}\operatorname{\mathscr{D}}_{\psi}\!\big{(}{\boldsymbol{\mu}}_{i}{+}{\boldsymbol{\epsilon}}\big{)};\,{\boldsymbol{\mu}}_{i}{=}\operatorname{\mathscr{E}}_{\phi}\!\big{(}\hat{S}_{i}\big{)}\!\right\},$ (4)
@@
Bound computation. From the PAC-Bayesian Theorem 2.1, using a prior $P_{\mathbf{0}}={\mathcal{N}}({\mathbf{0}},{\mathbf{I}})$ over the latent representation space and the comparator function of Equation (1), we obtain the following upper bound on the expected loss:
$b^{*}=\operatorname*{\mathrm{argsup}}_{b\in[0,1]}\quantity{\mathrm{kl}\quantity(\operatorname*{\mathbb{E}}\operatorname{\widehat{\mathcal{L}}}_{S^{\prime}}(h_{\gamma^{\prime}}),b)\leq\frac{\tfrac{1}{2}\|{\boldsymbol{\mu}}\|^{2}+\ln\tfrac{2\sqrt{{|S^{\prime}|}}}{\delta}}{|S^{\prime}|}}.$ (5)
That is, with probability at least $1{-}\delta$, we have $\operatorname*{\mathbb{E}}{\mathcal{L}}_{D^{\prime}}(h_{\gamma^{\prime}}){\leq}b^{*}$, where the expectation comes from the stochastic latent space.
@@ 3.2 Sample Compression Hypernetworks
    (optionally) A message ${\boldsymbol{\omega}}\in\{-1,1\}^{b}=\Omega$ (a binary-valued vector of of size $b$).
@@
$\min_{\psi}\left\{\frac{1}{m-|{\mathbf{j}}|}\sum_{({\mathbf{x}},y)\in S_{\bar{\mathbf{j}}}}\ell\big{(}h_{\gamma}({\mathbf{x}}),y\big{)}\ \Bigg{|}\ \gamma=\operatorname{\mathscr{R}}_{\psi}(S_{{\mathbf{j}}},{\boldsymbol{\omega}})\right\}.$ (6)
@@
Training objective. Our goal is to learn parameters $\psi$, ${\phi_{1}}$ and ${\phi_{2}}$ such that, for any task ${\mathcal{D}}^{\prime}\sim{\mathbf{D}}$ producing $S^{\prime}\sim{\mathcal{D}}^{\prime}$, the resulting output gives rise to a downstream predictor $h_{\gamma^{\prime}}$ of low generalization loss ${\mathcal{L}}_{{\mathcal{D}}^{\prime}}(h_{\gamma^{\prime}}$), with
@@
Note that the learning objective is a surrogate for Equation (6), as the complement of the compression set $\left(S_{\bar{\mathbf{j}}}\right)$ is replaced by the query set $\hat{T}_{i}$ in Equation (8). The pseudocode of the proposed approach can be found in Appendix G.
Figure 2: The Sample Compression hypernetwork.
Bound computation. When the 0-1 loss is used, the generalization bound from Theorem 2.2 is computed, using a fixed size $c$ for the compression sets; given a dataset size $m^{\prime}$, we use $J=\{\operatorname{\mathbf{j}}\in\mathscr{P}(\mathbf{m^{\prime}})~{}:~{}|\operatorname{\mathbf{j}}|=c\}$ and a uniform probability distribution over all distinct compression sets (sets that are not permutations of one another): $p=P_{J}(\operatorname{\mathbf{j}})=\binom{m^{\prime}}{c}^{-1}~{}\forall\operatorname{\mathbf{j}}\in J$. When using the optional message compressor module, we set a message size $b$ and a uniform distribution over all messages of size $b$: $P_{\Omega}({\boldsymbol{\omega}})=2^{-b}~{}\forall{\boldsymbol{\omega}}\in\{-1,1\}^{b}$, leading to the following upper bound on the loss:
$b^{*}=\operatorname*{\mathrm{argsup}}_{r\in[0,1]}\left\{\sum_{k=1}^{K}\binom{m^{\prime}-c}{k}r^{k}(1-r)^{m^{\prime}-c-k}\geq p\hskip 1.42262pt2^{-b}\hskip 1.42262pt\delta\right\},$
with $K=|\overline{\operatorname{\mathbf{j}}}|\widehat{\mathcal{L}}_{S_{\overline{\operatorname{\mathbf{j}}}}}\big{(}h_{\gamma^{\prime}}\big{)}\,$.
@@
$b^{*}=\operatorname*{\mathrm{argsup}}_{b\in[0,1]}\quantity{\mathrm{kl}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S^{\prime}}(h_{\gamma^{\prime}}),b)\leq\frac{1}{|S^{\prime}|-c}\ln\frac{2^{b+1}\sqrt{{|S^{\prime}|-c}}}{p\cdot\delta}}.$
That is, with probability at least $1-\delta$, we have ${\mathcal{L}}_{{\mathcal{D}}^{\prime}}(h_{\gamma^{\prime}})\leq b^{*}$.
@@ 3.3 PAC-Bayes Sample Compression Hypernetworks
Recall that Theorem 2.4 is obtained by handling the message of the sample compress framework in a PAC-Bayesian fashion, enabling the use of a posterior distribution over continous messages. Hence, this motivates revisiting the sample compress hypernetwork of Section 3.2 by replacing the message compressor $\operatorname{\mathscr{M}}_{\phi_{2}}$ (outputting a binary vector) by the PAC-Bayes encoder of Section 3.1. We denote the latter $\operatorname{\mathscr{E}}_{\phi_{2}}$, whose task is to output the mean ${\boldsymbol{\mu}}\in{\mathbb{R}}^{b}$ of a posterior distribution $Q_{\Omega,{\boldsymbol{\mu}}}={\mathcal{N}}({\boldsymbol{\mu}},{\mathbf{I}})$ over a real-valued message space $\Omega={\mathbb{R}}^{b}$. The sample compressor remaining unchanged from Section 3.2, the PAC-Bayes Sample Compress architecture is expressed by $\operatorname{\mathscr{H}}_{\theta}(\cdot)=\operatorname{\mathscr{R}}_{\psi}(\operatorname{\mathscr{C}}_{\phi_{1}}(\cdot),\operatorname{\mathscr{E}}_{\phi_{2}}(\cdot))$.
Training objective. Based on the above formulation, the training objective of the proposed PAC-Bayes Sample Compression hypernetwork is
$\min_{\psi,{\phi_{1}},{\phi_{2}}}\left\{\frac{1}{n}\sum_{i=1}^{n}\operatorname*{\mathbb{E}}\widehat{\mathcal{L}}_{\hat{T}_{i}}(h_{\gamma_{i}})\,\Bigg{|}\,\gamma_{i}{=}\operatorname{\mathscr{R}}_{\psi}\!\big{(}\operatorname{\mathscr{C}}_{\phi_{1}}\big{(}\hat{S}_{i}\big{)},\operatorname{\mathscr{E}}_{\phi_{2}}\big{(}\hat{S}_{i}\big{)}{+}{\boldsymbol{\epsilon}}\big{)}\!\right\},$ (9)
@@
Bound computation. From Theorem 2.4, using a fixed compression set size $c$, a prior $P_{\Omega,{\mathbf{0}}}={\mathcal{N}}({\mathbf{0}},{\mathbf{I}})$ over the real-valued message space of size $b$, a uniform probability over the compression set choice $p=P_{J}(\operatorname{\mathbf{j}})=\binom{m^{\prime}}{c}^{-1}~{}\forall\operatorname{\mathbf{j}}\in J$, and the comparator function of Equation (1), we obtain the following upper bound on the expected loss:
$b^{*}=\operatorname*{\mathrm{argsup}}_{b\in[0,1]}\quantity{\mathrm{kl}\quantity(\operatorname*{\mathbb{E}}\operatorname{\widehat{\mathcal{L}}}_{S^{\prime}}(h_{\gamma^{\prime}}),b)\leq\frac{\tfrac{1}{2}\|{\boldsymbol{\mu}^{\prime}}\|^{2}+\ln\tfrac{2\sqrt{{|S^{\prime}|-c}}}{p\cdot\delta}}{|S^{\prime}|-c}}.$
That is, with probability at least $1-\delta$, we have $\operatorname*{\mathbb{E}}{\mathcal{L}}_{D^{\prime}}(h_{\gamma^{\prime}})\leq b^{*}$, where the expectation comes from the stochastic message treatment.
To compute the disintegrated bound of Theorem 2.5, we choose $\zeta$ to output the compression set $S_{\operatorname{\mathbf{j}}}={\mathcal{C}}_{{\phi_{1}}}\quantity(S^{\prime})$ and a normal distribution centered on the message $\operatorname{\mathcal{E}}_{{\phi_{2}}}(S^{\prime})$, denoted $Q_{\Omega}^{S}={\mathcal{N}}(\operatorname{\mathscr{E}}_{\phi_{2}}\quantity(S^{\prime}),{\mathbf{I}})$. Thus, we have $\zeta(S,P_{\Omega})=\quantity(S_{\operatorname{\mathbf{j}}},Q_{\Omega}^{S}).$ After sampling a message $\omega\sim Q_{\Omega}^{S}$, we define $\gamma^{\prime}={\mathcal{R}}_{\psi}\quantity(S_{\operatorname{\mathbf{j}}},\omega)$. Applying Theorem 2.5 with $\alpha=2$, we have the following upper bound on the loss :
$b^{*}=\operatorname*{\mathrm{argsup}}_{b\in[0,1]}\quantity{\mathrm{kl}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S^{\prime}}(h_{\gamma^{\prime}}),b)\leq\frac{\|{\boldsymbol{\mu}^{\prime}}\|^{2}+\ln\frac{16\sqrt{|S^{\prime}|-c}}{p\cdot\delta^{3}}}{|S^{\prime}|-c}}.$
That is, with probability at least $1-\delta$, we have $\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}^{\prime}}(h_{\gamma^{\prime}})\leq b^{*}.$
@@ 4.1 Implementation Details
In each task, we split our training datasets into train and validation datasets; for each meta-learning hypernetwork, the hyperparameters are selected according to the error on made on the validation datasets. Detailed hyperparameters used for each experiment are given in Appendix H.
DeepSet dataset encoding. The hypernetwork $\operatorname{\mathscr{H}}_{\theta}(S)$ must be invariant to the permutation of its input $S$: the order of the examples in the input dataset should not affect the resulting encoding. Modules such as FSPool (Zhang et al., 2020) or a transformer (Vaswani et al., 2017) ensure such property. Our experiments use a simpler mechanism that is inspired by the DeepSet method (Zaheer et al., 2017).
@@
For $\kappa$-class tasks, where $\kappa>2$: given a data-matrix ${\mathbf{X}}\in{\mathbb{R}}^{m\times d}$ and a one-hot encoding of the label ${\mathbf{Y}}$, the output of a DeepSet module is the embedding ${\mathbf{z}}\in{\mathbb{R}}^{d^{\prime}}$, obtained by first applying a fully-connected neural network $g_{\omega}:{\mathbb{R}}^{d}\to{\mathbb{R}}^{d^{\prime}+\kappa}$ to each row of ${\mathbf{X}}$, where the label representation have been appended, to obtain a matrix ${\mathbf{M}}\in{\mathbb{R}}^{m\times(d^{\prime}+\kappa)}$ and then aggregating the result column-wise: ${\mathbf{z}}=\frac{1}{m}{\mathbf{M}}^{T}\mathbf{1}$.
PAC-Bayes encoder / message compressor. The PAC-Bayes encoder takes as input a dataset and outputs a continuous representation. [^1] In a slight language abuse, from now on, this continuous representation will be refereed to as message, just like the output of the message compressor. It is composed of a DeepSet module, followed by a feedforward network. Its last activation function is a Tanh function, so that the message is close to ${\mathbf{0}}$, leading to a bound that is not penalized much from having an important representation size (see Equation 5).

### v3

@@ 3.1 PAC-Bayes Encoder-Decoder Hypernetworks
We depart from previous works on PAC-Bayesian meta-learning and from classical PAC-Bayes, which consider a hierarchy of (meta-)prior and (meta-)posterior distributions (e.g., Pentina & Lampert, 2014). Instead, we consider distributions over a latent representation learned by the hypernetwork $\operatorname{\mathscr{H}}_{\theta}$. To encourage this representation to focus on relevant information, we adopt an encoder-decoder architecture $\operatorname{\mathscr{H}}_{\theta}(\cdot)=\operatorname{\mathscr{D}}_{\psi}(\operatorname{\mathscr{E}}_{\phi}(\cdot))$, with $\theta=\{\phi,\psi\}$. The encoder $\operatorname{\mathscr{E}}_{\phi}$ compresses the relevant dataset information into a vector ${\boldsymbol{\mu}}\in{\mathbb{R}}^{|{\boldsymbol{\mu}}|}$ (typically, $|{\boldsymbol{\mu}}|{\ll}|\gamma|$). The vector ${\boldsymbol{\mu}}$ is then treated as the mean of the Gaussian posterior distribution $Q_{\boldsymbol{\mu}}{\,=\,}{\mathcal{N}}({\boldsymbol{\mu}},{\mathbf{I}})$ over the latent representation space, such that the decoder $\operatorname{\mathscr{D}}_{\psi}$ maps the realizations drawn from $Q_{\boldsymbol{\mu}}$ to the downstream parameters $\gamma$. Figure 2 illustrates this architecture.
Figure 2: The PAC-Bayes hypernetwork.
Training objective. Based on the above, learning the proposed encoder-decoder hypernetwork amount to solve
$\min_{\psi,\phi}\left\{\frac{1}{n}\sum_{i=1}^{n}\operatorname*{\mathbb{E}}\widehat{\mathcal{L}}_{\hat{T}_{i}}(h_{\gamma_{i}})\,\Bigg{|}\,\gamma_{i}{=}\operatorname{\mathscr{D}}_{\psi}\!\big{(}{\boldsymbol{\mu}}_{i}{+}{\boldsymbol{\epsilon}}\big{)};\,{\boldsymbol{\mu}}_{i}{=}\operatorname{\mathscr{E}}_{\phi}\!\big{(}\hat{S}_{i}\big{)}\!\right\},$
@@
Bound computation.  Given a new task sample $S^{\prime}\,{\sim}\,({\mathcal{D}}^{\prime})^{m^{\prime}}$, we obtain from the PAC-Bayesian Theorem 2.1 (using a prior $P_{\mathbf{0}}={\mathcal{N}}({\mathbf{0}},{\mathbf{I}})$ over the latent representation space and the comparator function of Equation (1)), the following upper bound on the expected loss according to ${\mathcal{D}}^{\prime}$:
$\tau^{*}\!=\!\operatorname*{\mathrm{argsup}}_{\tau\in[0,1]}\!\quantity{\!\mathrm{kl}\quantity(\operatorname*{\mathbb{E}}\operatorname{\widehat{\mathcal{L}}}_{S^{\prime}}(h_{\gamma^{\prime}}),\tau\!)\!\leq\!\frac{\tfrac{1}{2}\|{\boldsymbol{\mu}}\|^{2}\!+\!\ln\!\tfrac{2\sqrt{{m^{\prime}}}}{\delta}}{m^{\prime}}\!},\!$ (4)
where ${\boldsymbol{\mu}}$ is the mean of the Gaussian posterior distribution. That is, with probability at least $1{-}\delta$, we have $\operatorname*{\mathbb{E}}{\mathcal{L}}_{{\mathcal{D}}^{\prime}}(h_{\gamma^{\prime}})\,{\leq}\,\tau^{*}$, where the expectation comes from the stochastic latent space.
@@ 3.2 Sample Compression Hypernetworks
    (optionally) A message ${\boldsymbol{\omega}}\in\{-1,1\}^{b}=\Omega$, that is a binary-valued vector of size $b$.
@@
$\min_{\psi}\!\Bigg{\{}\frac{1}{m-|{\mathbf{j}}|}\sum_{({\mathbf{x}},y)\in S_{\bar{\mathbf{j}}}}\!\!\!\ell\big{(}h_{\gamma}({\mathbf{x}}),y\big{)}\ \Bigg{|}\ \gamma=\operatorname{\mathscr{R}}_{\psi}(S_{{\mathbf{j}}},{\boldsymbol{\omega}})\Bigg{\}}.\!$ (5)
@@
Figure 3: The Sample Compression hypernetwork.
Training objective.  Our goal is to learn parameters $\psi$, ${\phi_{1}}$ and ${\phi_{2}}$ such that, for any task ${\mathcal{D}}^{\prime}{\sim}\,{\mathbf{D}}$ producing $S^{\prime}{\sim}\,({\mathcal{D}}^{\prime})^{m^{\prime}}$, the resulting output gives rise to a downstream predictor $h_{\gamma^{\prime}}$ of low generalization loss ${\mathcal{L}}_{{\mathcal{D}}^{\prime}}(h_{\gamma^{\prime}}$), with
@@
Note that the learning objective is a surrogate for Equation (5), as the complement of the compression set $S_{\bar{\mathbf{j}}}$ is replaced by the query set $\hat{T}_{i}$ in Equation (7). The pseudocode of the proposed approach can be found in Appendix G.
Bound computation.  When the 0-1 loss is used, the generalization bound from Theorem 2.2 is computed, using a fixed size $c$ for the compression sets; given a dataset size $m^{\prime}$, we use $J=\{\operatorname{\mathbf{j}}\in\mathscr{P}(\mathbf{m^{\prime}})~{}:~{}|\operatorname{\mathbf{j}}|=c\}$ and a uniform probability distribution over all distinct compression sets (sets that are not permutations of one another): $p=P_{J}(\operatorname{\mathbf{j}})=\binom{m^{\prime}}{c}^{{}_{-1}}~{}\forall\operatorname{\mathbf{j}}\in J$. When using the optional message compressor module, we set a message size $b$ and a uniform distribution over all messages of size $b$: $P_{\Omega}({\boldsymbol{\omega}})=2^{-b}~{}\forall{\boldsymbol{\omega}}\in\{-1,1\}^{b}$, leading to the following upper bound on the loss:
$\tau^{*}=\operatorname*{\mathrm{argsup}}_{\tau\in[0,1]}\left\{\sum_{k=1}^{K}\binom{m^{\prime}-c}{k}\tau^{k}(1-\tau)^{m^{\prime}-c-k}\geq p\hskip 1.42262pt2^{-b}\hskip 1.42262pt\delta\right\},$
with $K=|\overline{\operatorname{\mathbf{j}}}|\widehat{\mathcal{L}}_{S^{\prime}_{\overline{\operatorname{\mathbf{j}}}}}\big{(}h_{\gamma^{\prime}}\big{)}\,$.
@@
$\tau^{*}=\operatorname*{\mathrm{argsup}}_{\tau\in[0,1]}\quantity{\mathrm{kl}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S^{\prime}_{\overline{\operatorname{\mathbf{j}}}}}(h_{\gamma^{\prime}}),\tau)\leq\frac{1}{m^{\prime}-c}\ln\frac{2^{b+1}\sqrt{{m^{\prime}-c}}}{p\,\delta}}.$
That is, ${\mathcal{L}}_{{\mathcal{D}}^{\prime}}(h_{\gamma^{\prime}})\leq\tau^{*}$ with probability at least $1-\delta$.
@@ 3.3 PAC-Bayes Sample Compression Hypernetworks
Recall that Theorem 2.4 is obtained by handling the message of the sample compress framework in a PAC-Bayesian fashion, enabling the use of a posterior distribution over continous messages. Hence, this motivates revisiting the sample compress hypernetwork of Section 3.2 by replacing the message compressor $\operatorname{\mathscr{M}}_{\phi_{2}}$ (outputting a binary vector) by the PAC-Bayes encoder of Section 3.1. We denote the latter $\operatorname{\mathscr{E}}_{\phi_{2}}$, whose task is to output the mean ${\boldsymbol{\mu}}\in{\mathbb{R}}^{b}$ of a posterior distribution $Q_{\Omega,{\boldsymbol{\mu}}}={\mathcal{N}}({\boldsymbol{\mu}},{\mathbf{I}})$ over a real-valued message space $\Omega={\mathbb{R}}^{b}$. The sample compressor remaining unchanged from Section 3.2, the PAC-Bayes Sample Compress architecture is expressed by $\operatorname{\mathscr{H}}_{\theta}(\cdot)=\operatorname{\mathscr{R}}_{\psi}(\operatorname{\mathscr{C}}_{\phi_{1}}(\cdot),\operatorname{\mathscr{E}}_{\phi_{2}}(\cdot))$. Figure 4 illustrates the resulting architecture.
Training objective.  Based on the above formulation, we obtain the following training objective:
$\min_{\psi,{\phi_{1}},{\phi_{2}}}\!\!\left\{\frac{1}{n}\sum_{i=1}^{n}\operatorname*{\mathbb{E}}\widehat{\mathcal{L}}_{\hat{T}_{i}}(h_{\gamma_{i}})\,\Bigg{|}\,\gamma_{i}{=}\operatorname{\mathscr{R}}_{\psi}\!\big{(}\operatorname{\mathscr{C}}_{\phi_{1}}\big{(}\hat{S}_{i}\big{)},\operatorname{\mathscr{E}}_{\phi_{2}}\big{(}\hat{S}_{i}\big{)}{+}{\boldsymbol{\epsilon}}\big{)}\!\!\right\}\!,$
@@
Bound computation.  From Theorem 2.4, using a fixed compression set size $c$, a prior $P_{\Omega,{\mathbf{0}}}={\mathcal{N}}({\mathbf{0}},{\mathbf{I}})$ over the real-valued message space of size $b$, a uniform probability over the compression set choice $p=P_{J}(\operatorname{\mathbf{j}})=\binom{m^{\prime}}{c}^{{}_{-1}}~{}\forall\operatorname{\mathbf{j}}\in J$, and the comparator function of Equation (1), we obtain the following upper bound on the expected loss:
$\tau^{*}=\operatorname*{\mathrm{argsup}}_{\tau\in[0,1]}\quantity{\mathrm{kl}\quantity(\operatorname*{\mathbb{E}}\operatorname{\widehat{\mathcal{L}}}_{S^{\prime}_{\overline{\operatorname{\mathbf{j}}}}}(h_{\gamma^{\prime}}),\tau)\leq\frac{\tfrac{1}{2}\|{\boldsymbol{\mu}^{\prime}}\|^{2}+\ln\tfrac{2\sqrt{{m^{\prime}-c}}}{p\cdot\delta}}{m^{\prime}-c}}.$
That is, with probability at least $1-\delta$, we have $\operatorname*{\mathbb{E}}{\mathcal{L}}_{D^{\prime}}(h_{\gamma^{\prime}})\leq\tau^{*}$, where the expectation comes from the stochastic message treatment.
To compute the disintegrated bound of Theorem 2.5, we choose $\zeta$ to output the compression set $S^{\prime}_{\operatorname{\mathbf{j}}}={\mathcal{C}}_{{\phi_{1}}}(S^{\prime})$ and a normal distribution centered on the message $\operatorname{\mathcal{E}}_{{\phi_{2}}}(S^{\prime})$, denoted $Q_{\Omega}^{S}={\mathcal{N}}(\operatorname{\mathscr{E}}_{\phi_{2}}(S^{\prime}),{\mathbf{I}})$. Thus, we have $\zeta(S^{\prime},P_{\Omega})=(S^{\prime}_{\operatorname{\mathbf{j}}},Q_{\Omega}^{S}).$ After sampling a message $\omega\sim Q_{\Omega}^{S}$, we define $\gamma^{\prime}={\mathcal{R}}_{\psi}(S^{\prime}_{\operatorname{\mathbf{j}}},\omega)$. Applying Theorem 2.5 with $\alpha=2$, we have the following upper bound on the loss :
$\tau^{*}=\operatorname*{\mathrm{argsup}}_{\tau\in[0,1]}\quantity{\mathrm{kl}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S^{\prime}_{\overline{\operatorname{\mathbf{j}}}}}(h_{\gamma^{\prime}}),\tau)\leq\frac{\|{\boldsymbol{\mu}^{\prime}}\|^{2}+\ln\frac{16\sqrt{m^{\prime}-c}}{p\cdot\delta^{3}}}{m^{\prime}-c}}.$
That is, $\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}^{\prime}}(h_{\gamma^{\prime}})\leq\tau^{*}$ with probability at least $1-\delta$.
Figure 4: The PAC-Bayes Sample Compression hypernetwork.
@@ 4.1 Implementation Details
In each task, we split our training datasets into train and validation datasets; for each meta-learning hypernetwork, the hyperparameters are selected according to the error made on the validation datasets. Detailed hyperparameters used for each experiment are given in Appendix H. [^1] Our code is available at https://github.com/GRAAL-Research/DeepRM.
DeepSet dataset encoding.  The hypernetwork $\operatorname{\mathscr{H}}_{\theta}(S)$ must be invariant to the permutation of its input $S$: the order of the examples in the input dataset should not affect the resulting encoding. Modules such as FSPool (Zhang et al., 2020) or a transformer (Vaswani et al., 2017) ensure such property. Our experiments use a simpler mechanism that is inspired by the DeepSet module (Zaheer et al., 2017).
Figure 5: Illustration of decision boundaries generated by downstream predictors $h_{\gamma}$, when parameters $\gamma$ are given by a PAC-Bayes decoder $\operatorname{\mathscr{D}}_{\psi}({\boldsymbol{\mu}})$ with input ${\boldsymbol{\mu}}=(\mu_{1},\mu_{2})\in{\mathbb{R}}^{2}$. The encoder generated the message $(0.97,-0.27)$ on shown training datapoints. Left: The first latent dimension is fixed ($\mu_{1}=0.97$), while the second varies ($\mu_{2}\in[-1,1]$). Right: The first dimension varies ($\mu_{1}\in[-1,1]$), while the second is fixed $\mu_{2}=-0.27$.
@@
[…]
PAC-Bayes encoder / message compressor. The PAC-Bayes encoder takes as input a dataset and outputs a continuous representation. [^2] In a slight language abuse, from now on, this continuous representation will be referred to as message, just like the output of the message compressor. It is composed of a DeepSet module, followed by a feedforward network. Its last activation function is a Tanh function, so that the message is close to ${\mathbf{0}}$, leading to a bound that is not penalized much from having an important representation size (see Equation (missing) 4).